# MLP Memory: A Retriever-Pretrained Memory for Large Language Models

### v2

@@ 5.3 Hallucination Reduction
[…]
We assess MLP Memory’s ability to reduce hallucinations using HaluEval (Li et al., 2023) across three generation tasks: dialogue, question answering, and summarization, where models must identify factual inconsistencies in generated content. As shown in Table 3, parametric methods (CPT and LoRA) severely degrade performance, confirming the risks of weight modification. MLP Memory consistently improves hallucination detection across all three domains, with gains of 9.68, 10.08, and 2.14 points respectively. These substantial improvements indicate that the retrieval patterns encoded in MLP Memory significantly help the model better distinguish factual from hallucinated content. The effectiveness across diverse generation contexts—from free-form dialogue to constrained summarization—suggests that MLP Memory’s learned retrieval behavior provides a general mechanism for grounding language generation in factual knowledge. This hallucination reduction, combined with strong QA performance and enhanced general capabilities, validates our core hypothesis that decoupling memorization from reasoning through retriever-pretrained external memory can enhance factual accuracy without the typical trade-offs of parametric or non-parametric approaches.
@@ 5.4 Ablation study
Ablation Setup We conduct ablation experiments across three GPT2 (Radford et al., 2019) scales: small (12 layers), medium (24 layers), and large (36 layers), paired with corresponding MLP Memory modules of 117M, 345M, and 774M parameters respectively. All experiments are evaluated on WikiText-103 (Merity et al., 2016) to investigate loss weighting and optimal layer selection.
@@
Which Layer Provides the Best Representation for MLP Memory? While NN-LM performs best using the input to the final feedforward layer as the retrieval key, our MLP Memory consistently achieves optimal performance at around 70% of network depth, regardless of model scale. Our finding aligns with Memorizing Transformers (Burtsev et al., 2021), which also selected around 75% depth for optimal retrieval performance. We evaluate GPT2-small (12 layers), GPT2-medium (24 layers), and GPT2-large (36 layers), attaching the MLP Memory to various transformer blocks. As shown in Figure 5(b), the x-axis indicates relative depth (20%–100%), and the y-axis shows min-max normalized perplexity (0% = best, 100% = worst). This consistent pattern across all model sizes contrasts with the NN-LM convention of using final-layer representations.
@@ 6 Related Work
Retrieval-Augmented Generation RAG (Lewis et al., 2021; Peng et al., 2023; Gao et al., 2022) mitigates hallucinations by grounding generation in external knowledge. Despite improving factual accuracy, RAG faces limitations: retrieval latency, coarse granularity, and limited LLM integration (Zhang et al., 2024). Recent work (Su et al., 2025) explores enhanced retrieval with LLM priors. Our approach proposes a parametric memory mimicking non-parametric retrieval, eliminating explicit document retrieval while preserving knowledge augmentation.
Memory-Augmented Language Models Various architectures explored memory augmentation, from Memory Networks (Weston et al., 2015) with explicit read-write components to Memory Transformers (Burtsev et al., 2021) with extended attention. LongMem (Wang et al., 2023) and MemoRAG (Qian et al., 2025) introduced decoupled architectures for long-term history storage. While these focus on context extension, our MLP memory expands across the entire pre-training corpus, enabling long-term generalizable knowledge storage.
MLP Architectures All-MLP architectures emerged as transformer alternatives, with gMLP (Liu et al., 2021) matching transformer performance and sparse MLPs (Yu et al., 2022) showing superior training efficiency. Studies (Geva et al., 2020) identified FFN layers as key-value memories in LLMs. Inspired by this, we propose pretraining an all-MLP memory as a non-parametric retriever, leveraging MLPs’ memorization capabilities for a compact, differentiable knowledge store.
@@ 7 Conclusion
The key advantage of MLP Memory lies in its efficiency and effectiveness. Our approach achieves 12.3% relative improvement on question-answering benchmarks, 5.2 points gain on general NLP tasks, and up to 10 points reduction in hallucinations—while delivering 2.5 faster inference than RAG and maintaining constant speed regardless of corpus size. Unlike parametric fine-tuning that risks catastrophic forgetting or non-parametric RAG that suffers from high latency, MLP Memory enhances model capabilities without these typical trade-offs.
@@
Appendix A The Use of Large Language Models
We utilized ChatGPT as an editorial assistant to enhance the manuscript’s language quality, correct grammatical errors, and ensure academic writing standards. All scientific contributions, including the research design, experimental procedures, and analytical interpretations, were originated and performed independently by the authors.
Appendix B Reproducibility Statements
Appendix C Implementation Details
@@
For the general NLP tasks in Section 5.2, we utilize a heterogeneous corpus constructed by aggregating several publicly available sources that cover diverse domains relevant to common NLP tasks. Following the methodology from (Geng et al., 2024), this corpus comprises WikiText-103 Merity et al. (2016) for encyclopedic content, Amazon Reviews He & McAuley (2016) for user-generated product feedback, CC-NEWS Hamborg et al. (2017) for journalistic content, and IMDB Maas et al. (2011) for movie reviews and discussions.
@@
For question answering benchmarks, following Cheng et al. (2024), we evaluate three Open Domain Question Answering datasets and HotpotQA using the Exact Match (EM) metric. For long-form QA evaluation, we employ three complementary metrics: MC1, which measures whether the model assigns the highest likelihood to the most accurate answer; MC2, which sums the normalized probabilities over all correct answers; and MC3, which evaluates whether the model assigns a higher average likelihood to true answers than to false ones. We report the average of these three metrics as the final performance measure for long-form QA tasks. For general NLP tasks, following the methodology from Shi et al. (2022b), we report results using the domain-conditional PMI scoring rule Holtzman et al. (2021). For hallucination reduction evaluation, we use accuracy as the primary metric to assess the model’s ability to generate factually correct responses.
@@
Appendix D Sensitivity to Interpolation Weight
@@
Appendix E Inference Efficiency Analysis
@@
Appendix F Scaling law
Setup
We conduct scaling law experiments using standard decoder-only models and our overall model architecture. As baselines, we use four GPT-2 Radford et al. (2019) variants with increasing parameter counts: GPT2-small (124M), GPT2-medium (345M), GPT2-large (774M), and GPT2-xl (1.5B). For MLP Memory, we define three configurations: small (124M), medium (335M), and large (774M) that align with the scaling trend of standard architectures. The MLP Memory module is externally integrated with a matching-sized GPT-2 variant, resulting in total parameter counts of approximately 248M, 710M, and 1.5B for the small, medium, and large configurations, respectively. All models are trained on two datasets: WikiText-103 Merity et al. (2016) (around 100M tokens) and a mixed Web dataset (around 600M tokens). Following Shi et al. (2022a), our Web dataset combines diverse knowledge sources relevant to common NLP tasks, including WikiText-103, Amazon Reviews He & McAuley (2016), CC-NEWS Hamborg et al. (2017), and IMDB Maas et al. (2011).
Scaling law with model parameters
Following Kaplan et al. (2020), we model perplexity scaling as . Under fixed compute, we compare our architecture to continued GPT-2 training on WikiText-103 and Web datasets in terms of test perplexity scaling with model size . Results in Figure6 show our architecture demonstrates a steeper scaling curve than the decoder-only model, indicating improved scaling efficiency. The power-law scaling laws on WikiText-103 can be expressed as:
Scaling law with training compute
Appendix G Comparing Output Distribution Characteristics of LM, NN, and MLP Memory
@@
Appendix H Effect of Different NN Target Distributions
@@
Appendix I Sensitivity to in Target Distribution Generation
@@
Appendix J Case Study

### v3

@@ 5.4 Hallucination Reduction
| +MLP Mem | 66.869.68 | 64.0710.08 | 52.412.14 |
We assess MLP Memory’s ability to reduce hallucinations using HaluEval [Li et al., 2023] across three generation tasks: dialogue, question answering, and summarization, where models must identify factual inconsistencies in generated content. As shown in Table 3, parametric methods (CPT and LoRA) severely degrade performance, confirming the risks of weight modification. MLP Memory consistently improves hallucination detection across all three domains, with gains of 9.68, 10.08, and 2.14 points respectively. These substantial improvements indicate that the retrieval patterns encoded in MLP Memory significantly help the model better distinguish factual from hallucinated content. The effectiveness across diverse generation contexts—from free-form dialogue to constrained summarization—suggests that MLP Memory’s learned retrieval behavior provides a general mechanism for grounding language generation in factual knowledge. This hallucination reduction, combined with strong QA performance and enhanced general capabilities, validates our core hypothesis that decoupling memorization from reasoning through retriever-pretrained external memory can enhance factual accuracy without the typical trade-offs of parametric or non-parametric approaches.
@@ 5.5 Ablation study
Ablation Setup We conduct ablation experiments across three GPT2 [Radford et al., 2019] scales: small (12 layers), medium (24 layers), and large (36 layers), paired with corresponding MLP Memory modules of 117M, 345M, and 774M parameters respectively. All experiments are evaluated on WikiText-103 [Merity et al., 2016] to investigate loss weighting and optimal layer selection.
@@
Which Layer Provides the Best Representation for MLP Memory? While NN-LM performs best using the input to the final feedforward layer as the retrieval key, our MLP Memory consistently achieves optimal performance at around 70% of network depth, regardless of model scale. Our finding aligns with Memorizing Transformers [Burtsev et al., 2021], which also selected around 75% depth for optimal retrieval performance. We evaluate GPT2-small (12 layers), GPT2-medium (24 layers), and GPT2-large (36 layers), attaching the MLP Memory to various transformer blocks. As shown in Figure 6(b), the x-axis indicates relative depth (20%–100%), and the y-axis shows min-max normalized perplexity (0% = best, 100% = worst). This consistent pattern across all model sizes contrasts with the NN-LM convention of using final-layer representations.
@@ 6 Related Work
Retrieval-Augmented Generation RAG [Lewis et al., 2021, Peng et al., 2023, Gao et al., 2022] mitigates hallucinations by grounding generation in external knowledge. Despite improving factual accuracy, RAG faces limitations: retrieval latency, coarse granularity, and limited LLM integration [Zhang et al., 2024]. Recent work [Su et al., 2025] explores enhanced retrieval with LLM priors. Our approach proposes a parametric memory mimicking non-parametric retrieval, eliminating explicit document retrieval while preserving knowledge augmentation.
Memory-Augmented Language Models Various architectures explored memory augmentation, from Memory Networks [Weston et al., 2015] with explicit read-write components to Memory Transformers [Burtsev et al., 2021] with extended attention. LongMem [Wang et al., 2023] and MemoRAG [Qian et al., 2025] introduced decoupled architectures for long-term history storage. While these focus on context extension, our MLP memory expands across the entire pre-training corpus, enabling long-term generalizable knowledge storage.
MLP Architectures All-MLP architectures emerged as transformer alternatives, with gMLP [Liu et al., 2021] matching transformer performance and sparse MLPs [Yu et al., 2022] showing superior training efficiency. Studies [Geva et al., 2020] identified FFN layers as key-value memories in LLMs. Inspired by this, we propose pretraining an all-MLP memory as a non-parametric retriever, leveraging MLPs’ memorization capabilities for a compact, differentiable knowledge store.
@@ 7 Conclusion
The key advantage of MLP Memory lies in its efficiency and effectiveness, as our architecture exhibits stronger scaling behavior than decoder-only models. On downstream tasks, our approach achieves 12.3% relative improvement on question-answering benchmarks, 5.2 points gain on general NLP tasks, and up to 10 points reduction in hallucinations—while delivering 2.5 faster inference than RAG and maintaining constant speed regardless of corpus size. Unlike parametric fine-tuning that risks catastrophic forgetting or non-parametric RAG that suffers from high latency, MLP Memory enhances model capabilities without these typical trade-offs.
@@
Appendix A Implementation Details
@@
For the general NLP tasks in Section 5.3, we utilize a heterogeneous corpus constructed by aggregating several publicly available sources that cover diverse domains relevant to common NLP tasks. Following the methodology from [Geng et al., 2024], this corpus comprises WikiText-103 Merity et al. [2016] for encyclopedic content, Amazon Reviews He and McAuley [2016] for user-generated product feedback, CC-NEWS Hamborg et al. [2017] for journalistic content, and IMDB Maas et al. [2011] for movie reviews and discussions.
@@
For question answering benchmarks, following Cheng et al. [2024], we evaluate three Open Domain Question Answering datasets and HotpotQA using the Exact Match (EM) metric. For long-form QA evaluation, we employ three complementary metrics: MC1, which measures whether the model assigns the highest likelihood to the most accurate answer; MC2, which sums the normalized probabilities over all correct answers; and MC3, which evaluates whether the model assigns a higher average likelihood to true answers than to false ones. We report the average of these three metrics as the final performance measure for long-form QA tasks. For general NLP tasks, following the methodology from Shi et al. [2022b], we report results using the domain-conditional PMI scoring rule Holtzman et al. [2021]. For hallucination reduction evaluation, we use accuracy as the primary metric to assess the model’s ability to generate factually correct responses.
@@
Appendix B Sensitivity to Interpolation Weight
@@
Appendix C Inference Efficiency Analysis
@@
Appendix D Comparing Output Distribution Characteristics of LM, NN, and MLP Memory
@@
Appendix E Effect of Different NN Target Distributions
@@
Appendix F Sensitivity to in Target Distribution Generation
@@
Appendix G Case Study